OpenStudy (anonymous):

HELP! The sales tax in one state is 7%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a CD player with a selling price of $150. a. f(x) = x – 7; $148.93 c. f(x) = 7 + x; $151.07 b. f(x) = 0.07x; $10.50 d. f(x) =1.07x; $160.50

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 7}}\%\ of\ x \implies \cfrac{{\color{brown}{ 7}}}{100}\cdot x\implies {\color{brown}{ 7}}\div 100\cdot x \\ \quad \\ \textit{full price plus tax}\implies x+({\color{brown}{ 7}}\div 100\cdot x)\)